What is the energy of a photon (in joules) emitted during a transition from ni = (5.00x100) state to the nf = (2.00x100) state of the hydrogen atom?

1
Expert's answer
2021-10-07T03:45:10-0400

Rydberg formula, "\frac{1}{\lambda}=R_H(1/n_f^2-1/n_i^2)"

"n_f=2, n_i=5"

"R_H=10973731.56816"

Energy of a photon, "E=hf=hc/\lambda=hc\times R_H(1/n_f^2-1/n_i^2)"

"E=6.626\times 10^{-34}\times 3\times 10^8\times 10973731.56816(1/2^2-1/5^2)=4.581\times 10^{-19}J"